# LanEvil: Benchmarking the Robustness of Lane Detection to Environmental Illusions

### v1

@@ 1. Introduction
Lane detection aims to identify the location of lane lines or road edges (Pan et al., 2018; Qin et al., 2020; Zheng et al., 2021; Qin et al., 2022; Jin et al., 2023), which now serves as the foundation for many driving assistant functions in the real-world auto-driving vehicles (Hillel et al., 2014), such as lane centering and adaptive cruise control.
@@
Based on the pipeline, we rigorously analyze the critical influence factors in LD scenarios including Dynamic Objects, Static Facilities, and Environmental Conditions, and systematically design 14 illusion types with 5 severity levels that collectively address a wide spectrum of real-world environmental illusions on the road and propose the LanEvil LD robustness evaluation benchmark. Overall, our benchmark encompasses 94 cases with editable 3D environments and a 90,292 sampled image dataset with 40,000 clean training images and 50,292 test images. Leveraging LanEvil , we conducted extensive experiments to benchmark the robustness of commonly adopted LD models, where we observed severe performance degeneration (-5.37% Accuracy and -10.70% F1-Score on average). Notably, the shadow effect caused the most considerable reduction (-7.39% Accuracy). To improve model robustness against environmental illusions, we propose the Attention Area Mixing (AAM) approach, a novel noise defense baseline leveraging hard examples, achieving a 3.76% improvement over vanilla detectors towards these illusions. To better demonstrate the potential of our LanEvil, we conducted joint simulation experiments on commercial autonomous driving systems (OpenPilot (comma, 2017) and Apollo (Baidu, 2022)), where we observed incorrect decisions resulting in car accidents. Finally, we also conducted case studies on real-world scenarios containing environmental illusions, which demonstrate the threats in practice. We hope this paper will raise awareness regarding potential security threats in autonomous scenarios. Our contributions are as follows:
@@ 2.2. Lane Detection Datasets
Besides the above common datasets, the robustness benchmarks that study natural corruption in LD scenarios are rare. Some pioneering benchmarks are devoted to assessing the robustness of image classification (Hendrycks and Dietterich, 2019; Zhang et al., 2021; Guo et al., 2023; Liu et al., 2020a; Xiao et al., 2023a, b) and object detection (Hendrycks and Dietterich, 2019; Liu et al., 2023c, 2022, 2021a; Tang et al., 2021; Zhang et al., 2023; Jiang et al., 2023; Zhang et al., 2024a; Wang et al., 2024; Zhang et al., 2024b) against common perturbations such as blur, weather conditions, and digital corruption. Recently, some studies have proposed to investigate the perception robustness of autonomous driving against common corruption (Dong et al., 2023; Kong et al., 2023). However, these works primarily focus on general 3D perception tasks such as detection and segmentation, and the generated corruption (e.g., motion blur) is not specially designed for LD. Moreover, there also exist some studies that generate specially designed lane-like adversarial attacks for autonomous driving (Sato et al., 2021; Boloor et al., [n. d.]), which is out of the scope of this paper.
@@ 5.3. Results on Noise Defense Methods
In this section, we further investigated the effectiveness of our proposed AAM method and existing noise defense methods on our LanEvil dataset. Specifically, we choose ResNet-18 as the backbone for the LaneATT and UFLD models and apply PGD adversarial training (Madry et al., 2017), cutout (DeVries and Taylor, 2017), copy-paste (Ghiasi et al., 2021), Augment HSV and MixUP (Zhang et al., 2017) for the noise defense methods as they improve the model robustness towards adversarial noises (Liu et al., 2019, 2020b, 2023b, 2023a) or natural corruption.

### v4

@@ 1. Introduction
Lane detection aims to identify the location of lane lines or road edges (Pan et al., 2018; Qin et al., 2020; Zheng et al., 2021; Qin et al., 2022; Jin et al., 2023; liang2024object), which now serves as the foundation for many driving assistant functions in the real-world auto-driving vehicles (Hillel et al., 2014), such as lane centering and adaptive cruise control.
@@
Based on the pipeline, we rigorously analyze the critical influence factors in LD scenarios including Dynamic Objects, Static Facilities, and Environmental Conditions, and systematically design 14 illusion types with 5 severity levels that collectively address a wide spectrum of real-world environmental illusions on the road and propose the LanEvil LD robustness evaluation benchmark. Overall, our benchmark encompasses 94 cases with editable 3D environments and a 90,292 sampled image dataset with 40,000 clean training images and 50,292 test images. Leveraging LanEvil , we conducted extensive experiments to benchmark the robustness of commonly adopted LD models, where we observed severe performance degeneration (-5.37% Accuracy and -10.70% F1-Score on average). Notably, the shadow effect caused the most considerable reduction (-7.39% Accuracy). To improve model robustness against environmental illusions, we propose the Attention Area Mixing (AAM) approach, a novel noise defense baseline leveraging hard examples, achieving a 3.76% improvement over vanilla detectors towards these illusions. To better demonstrate the potential of our LanEvil, we conducted joint simulation experiments on commercial autonomous driving systems (OpenPilot (comma, 2017) and Apollo (Baidu, 2022)), where we observed incorrect decisions resulting in car accidents. Finally, we also conducted case studies on real-world scenarios containing environmental illusions, which demonstrate the threats in practice. We hope this paper will raise awareness regarding potential security threats (Liang et al., 2023b; Liu et al., 2023e, b; Liang et al., 2024a, b; Li et al., 2024) in autonomous scenarios. Our contributions are as follows:
@@ 2.2. Lane Detection Datasets
Besides the above common datasets, the robustness benchmarks that study natural corruption in LD scenarios are rare. Some pioneering benchmarks are devoted to assessing the robustness of image classification (Hendrycks and Dietterich, 2019; Zhang et al., 2021; Guo et al., 2023; Liu et al., 2020a; Xiao et al., 2023a, b) and object detection (Hendrycks and Dietterich, 2019; Liu et al., 2023d, 2022, 2021a; Tang et al., 2021; Zhang et al., 2023; Jiang et al., 2023; Zhang et al., 2024a; Wang et al., 2024; Zhang et al., 2024b; Liang et al., 2021, 2020; Wei et al., 2018; Liang et al., 2022b, a; Wang et al., 2023; Liu et al., 2023a; He et al., 2023b; Liu et al., 2023f; He et al., 2023a; Muxue et al., [n. d.]; Lou et al., 2024) against common perturbations such as blur, weather conditions, and digital corruption. Recently, some studies have proposed to investigate the perception robustness of autonomous driving against common corruption (Dong et al., 2023; Kong et al., 2023). However, these works primarily focus on general 3D perception tasks such as detection and segmentation, and the generated corruption (e.g., motion blur) is not specially designed for LD. Moreover, there also exist some studies that generate specially designed lane-like adversarial attacks for autonomous driving (Sato et al., 2021; Boloor et al., [n. d.]), which is out of the scope of this paper.
@@ 5.3. Results on Noise Defense Methods
In this section, we further investigated the effectiveness of our proposed AAM method and existing noise defense (Sun et al., 2023; Liu et al., 2023d; Liang et al., 2023a, 2024c) methods on our LanEvil dataset. Specifically, we choose ResNet-18 as the backbone for the LaneATT and UFLD models and apply PGD adversarial training (Madry et al., 2017), cutout (DeVries and Taylor, 2017), copy-paste (Ghiasi et al., 2021), Augment HSV and MixUP (Zhang et al., 2017) for the noise defense methods as they improve the model robustness towards adversarial noises (Liu et al., 2019, 2020b, 2023c, 2023a) or natural corruption.